OpenStudy (anonymous):

What is the graph of this parametric equation? x = 6 sin t, y = 6 cos t; the domain: 0 ≤ t ≤ 2pi

OpenStudy (amorfide):

you want to get the equations in terms of sint and cost x=6sint divide by 6 x/6=sint y=6cost divide by 6 y/6=cost then you know that \[\sin^{2}t+\cos^{2}t=1\] so you square them both, and add them together to get \[(\frac{ x }{ 6 })^{2} + (\frac{ y }{ 6 })^{2}= \sin^{2}t+\cos^{2}t\] \[(\frac{ x }{ 6 })^{2} + (\frac{ y }{ 6 })^{2} = 1\]

OpenStudy (anonymous):

i'm following you so far, just not sure on the graphing

OpenStudy (amorfide):

since the domain for t was from 0<t<2pi you know that it must be in all quadrants you know \[x^{2}+y^{2}=r^{2}\] is a circle centre (0,0) radius r since the x and y are divided by the same number, it will represent a circle you could expand out the brackets if you prefer, and then you can surely draw a circle graph

OpenStudy (anonymous):

oh okay thank you!! that really cleared it up for me i understand now :)

OpenStudy (amorfide):

no problem I am glad I could help